OpenStudy (anonymous):

Tracy ran for 2/3 hours and Reina ran for 1/6 hours. How many more hours did Tracy run? 1/4 1/2 1/3 1/6

OpenStudy (anonymous):

2/3-1/6 =4/6-1/6 =3/6 =1/2